An investor in the United States bought a one-year Brazilian security valued at 370,000 Brazilian reals (R$). The U.S. dollar equivalent was $320,000. The Brazilian security earned 12 percent during the year, but the Brazilian real depreciated 5 cents against the U.S. dollar during the time period ($.86 to $.81). After transferring the funds back to the United States, what was the investor’s return on her $320,000?
Answer #1

Investment in Brazilian Security = 370000 Brazilian

Brazilian security earned 12 percent during the year

Value after an year of Brazilian Security = 370000*1.12 = 414400 Brazilian

1 Brazilian = $0.81 after an year

So, 414400 Brazilian is 414400*0.81 = $335664

Initial Investment = $ 320000

Rate of Return = (335664/320000)-1 = 4.89%